The perimeter of a rectangle is 108 meters. The ratio of the width to the length is 4 to 5. Find the width. If necessary, round to the nearest hundredth
Answer by edjones(8007)

Let L=length, w=width
L/w=5/4
wL/w=5w/4
L=5w/4
.
2L+2w=P
(2*5w/4)+2w=108 Substitute 5w/4 for L
5w/2 + 2w=108
5w+4w=216 multiply each side by 2 to eliminate fractions.
9w=216
9w/9=216/9
w=24 m
.
L=5w/4
=5*24/4
=30 m
.
Check:
24*2 + 30*2 = 108
48+60=108 true
.
Ed